A social club charters a bus at a cost of $8700 to take a group of members on an excursion to Atlantic City. At the last minute, 2 people in the group decide not to go. This raises the transportation cost per person by $5. How many people originally intended to take the trip?
Answer by edjones(8007)

Let the number of people going on the trip initially be p.
(8700/p)+5=8700/(p-2)
8700(p-2)+5(p^2-2p)=8700p
8700p-17400+5p^2-10p=8700p Multiply each side by p(p-2)
5p^2-10p-17400=0
p^2-2p-3480=0
(p+58)(p-60)=0
p=60 persons
.
Ed